# Principal Component Flow Map Learning of PDEs from incomplete, limited, and noisy data

### v1

@@ Abstract
We present a computational technique for modeling the evolution of dynamical systems in a reduced basis, with a focus on the challenging problem of modeling partially-observed partial differential equations (PDEs) on high-dimensional non-uniform grids. We address limitations of previous work on data-driven flow map learning in the sense that we focus on noisy and limited data to move toward data collection scenarios in real-world applications. Leveraging recent work on modeling PDEs in modal and nodal spaces, we present a neural network structure that is suitable for PDE modeling with noisy and limited data available only on a subset of the state variables or computational domain. In particular, spatial grid-point measurements are reduced using a learned linear transformation, after which the dynamics are learned in this reduced basis before being transformed back out to the nodal space. This approach yields a drastically reduced parameterization of the neural network compared with previous flow map models for nodal space learning. This primarily allows for smaller training data sets, but also enables reduced training times.
@@ 1.1 Motivation
When modeling high-dimensional systems, e.g. systems modeled traditionally by partial differential equations (PDEs) with one or more spatial variables observed on a fine spatial grid, NN models require many parameters to account for different system behaviors. In moving from toy problems to real applications, a significant challenge throughout the field is the curse of dimensionality, which is the phenomenon that multiple system observations (training data samples) are required to accurately learn each model parameter without over-fitting. The more complex and high-dimensional a system is, the more parameters are needed in the model, and hence the more system observations are needed for training data. On the other hand, real-world data collection scenarios limit the spatial grid on which measurements are made, the time at which observations are made, the state variables that are actually observed, and the amount of data collected. If only a few noisy data are available to train a highly parameterized model, accuracy and generalizability suffer. Therefore, this paper addresses the immediate need for methods that require fewer parameters to enable efficient and scalable data-driven modeling of unknown high-dimensional dynamics while maintaining prediction accuracy and stability, especially when partial observations demand memory (i.e. a time history) of inputs. This paper focuses on learning flow maps to reduce the number of model parameters so that limited training data can be used productively.
@@ 1.2 Literature Review
Finally, two other papers, [15, 21], have similar goals of learning flows of dynamical systems in a reduced basis, albeit in different contexts. In [21], authors consider a partially-observed heat equation but observed at only 3 points throughout the domain, with little need for reduction of high-dimensional measurements and focusing more on reducing the dynamics. In [15], the authors consider a generalized Koopman framework similarly framed in the context of the Mori-Zwanzig formulation, but only consider a long ($10^{7}$ time steps), noiseless trajectory from the Lorenz ’96 ODE.
@@ 1.3 Our focus and contributions
Much like in [4], this paper considers data-driven flow map learning of PDEs from incomplete solution data. In other words, we consider the case of modeling an unknown PDE system using measurable data that represent partial information of the system. It is important to consider incomplete solution data because in real applications it is often difficult, if not impossible, to collect data of all the state variables of an unknown system in all areas of the relevant domain. It may even be difficult to identify all state variables in an unknown system.
Examples in [4] used up to $10,000$ noise-less training samples to train networks with up to hundreds of thousands of parameters with high accuracy. This demonstrated the not-so-surprising result that with nearly unlimited training time and data generated from densely packed randomized initial conditions, arbitrary accuracy is possible. However, the question now moves to the limited data scenario. As training data comes at a cost in applications, it is desirable to minimize the number of parameters in the model while still capturing accurate system dynamics. In what follows, we limit data to no more than 100 samples in all examples, and add noise. This combination of factors causes the nodal memory FML models of [4] to fail, motivating this exploration.
@@
Therefore, the main contribution of the paper is a memory-based framework for flow map learning of PDE in a learned lower-dimensional modal space from incomplete, limited, and noisy data. In particular, to reduce the measurements, we introduce unconstrained, constrained, and fixed (or pre-trained) approaches to learning common reduced bases for the training trajectories motivated by principal component analysis (PCA). To reduce the network parameterization, we derive the existence of an ODE flow map model and additionally propose a matrix form to further reduce the network parameterization. Finally, an adjustment to the overall ResNet structure radically improves the FML strategy’s robustness to noise. In a set of numerical examples, we compare these options against each other as well as the existing nodal model from [4] to demonstrate that our method is indeed highly effective and accurate.
@@ 3.2.1 Reduced basis for nodal measurements
$\delta_{n}^{(l)}=\left\|\mathbf{V}_{n}^{(l)}-\tilde{\mathbf{V}}_{n}\right\|^{2}_{2},$ (21)
for $l=1,\ldots,N_{traj}$ and $n=1,\ldots,N_{mem}+N_{rec}$ as in (17). Generally, if $N_{red}$ were fixed, we might seek an optimal basis of the form (19)
$\{b_{j}^{*}(\mathbf{x})\}_{j=1}^{N_{red}}=\arg\min_{\{b_{j}(\mathbf{x}\}_{j=1}^{N_{red}}}\sum_{l=1}^{N_{traj}}\sum_{n=1}^{N_{mem}+N_{rec}}\delta_{n}^{(l)}.$ (22)
that accurately approximates the solutions at all trajectories and time steps in the training data. We note that an optimal basis may be significantly different from those generating each training instance’s exact finite-term series representation of the finite-dimensional grid measurements.
@@
A focus of the ensuing technique is on finding matrices $\Pi_{in}$ and $\Pi_{out}$, that balance minimization of the error $\delta_{n}^{(l)}$ and the reduced dimension $N_{red}$ over the entire training dataset (17). As described in Section 3.5, several approaches are taken, motivated by a principal component analysis (PCA), also known as principal orthogonal decomposition (POD), approach. E.g., perhaps the most straightforward approach is to use fixed transformations based on the PCA weight matrix which can be pre-computed from the training data. This would enforce an accurate approximation of $\mathbf{V}_{n}$ by $\Pi_{out}\Pi_{in}\mathbf{V}_{n}$ along with $\Pi_{in}=\Pi_{out}^{T}$ and $\Pi_{out}$ having orthogonal columns.
@@ 3.2.2 Partial reduced dynamics
the solution of which has an approximate correspondence to the solution of the PDE in $\mathbf{V}$. This permits learning the evolution of the system in $\mathbb{R}^{N_{red}}$ instead of $\mathbb{R}^{N_{full}}$. When the governing PDE is unknown, the system for $\hat{\mathbf{V}}$ is unknown as well. Since $\mathbf{V}$ represent partial observations of the entire PDE system, $\hat{\mathbf{V}}$ represent partial observations of some entire system, say for $\tilde{\mathbf{U}}$, the modal coefficients of some approximation of the full system.
@@
$\bar{\mathbf{V}}_{n+1}=\Pi_{out}\Pi_{in}\bar{\mathbf{V}}_{n}+\Pi_{out}\mathcal{M}\left(\Pi_{in}\bar{\mathbf{V}}_{n-1},\ldots,\Pi_{in}\bar{\mathbf{V}}_{n-(N_{mem}-1)}\right),$ (29)
where $\mathcal{M}$ approximates the modal space coefficient flow map, and the bar notation denotes approximate quantities. After extensive numerical experimentation, we find that the transformation of $\bar{\mathbf{V}}_{n}$ into and out of the reduced space is essential to dealing with noisy nodal measurements. Essentially, $\Pi_{out}\Pi_{in}$ functions as a denoiser. If measurements are noise-free, it can be removed.
@@ 3.4 Determining reduced order dimension
Figure 3: Exs. 1, 2, 3 (left to right): Determining reduced order dimension.
Figure 3 shows the decay of the singular values in each of the three forthcoming numerical experiments with and without noise. We see, e.g., in Example 3 the noiseless singular values of the training data drop off gradually after $N_{red}=13$, while noise is dominant after $N_{red}=5$ in the noisy observations. Additionally, this SVD computation exposes potential basis elements $b_{j}(\mathbf{x})$ for $j=1,\ldots,N_{red}$ in the columns of $V_{red}$, which can be used to explain the decomposition and interpret the dynamics. Figures 4 and 5 show the 13- and 5-dimensional bases for the noiseless and noisy cases, respectively, in Example 3. We see similar shapes for the first 5 elements, while we also notice noise creeping into the $b_{4}(\mathbf{x})$ and $b_{5}(\mathbf{x})$ in the noisy measurements. This demonstrates why after this point, including extra basis elements is not advantageous as they will have a large noise component that may actually be detrimental to reconstructing the unknown dynamics.
Figure 4: Ex. 3 (noiseless): Reduced basis elements $b_{j}(\mathbf{x})$ for Fixed PC-FML.
Figure 5: Ex. 3 (noisy): Reduced basis elements $b_{j}(\mathbf{x})$ for Fixed PC-FML.
@@ 3.5 Model Training
Essentially, the options correspond to the amount of freedom we give the data and the network to learn the transformation. Indeed the inspiration for the general method was the Fixed regime, which can be relaxed to the Constrained regime’s orthogonality and symmetry or completely dropped in the Unconstrained regime. [^6] We admit that the network may find a better FML basis! Notably, in all cases the loss is measured in the nodal domain. This is significant, and differs from [2, 26], since the network parameters are trained to minimize loss in the nodal domain. If instead training is only executed in the modal domain, expanding back out could cause larger errors in the nodal domain, which is the focus.
@@ 4 Computational Studies
  • Time-step: fixed at $\Delta t=10^{-2}$.
@@
  • Recurrent loss: fixed at $N_{rec}=10$ steps.
@@
  • Noise: additive zero-mean Gaussian noise is considered at two noise levels, with standard deviations $\sigma=0$ and $\sigma=10^{-1}$, referred to generically as noiseless and noisy.
  • PC-FML network: networks $\mathbf{M}$ have $3$ hidden layers with either $10$ (Ex. 1) or $15$ (Exs. 2 and 3) nodes per layer. Hyperbolic tangent is used as the activation function.
  • Nodal FML network size: comparison networks from [4] have $5$ disassembly channels with $1$ hidden layer with either $N_{grid}$ (Exs. 1 and 2) or $N_{red}$ nodes per layer. The assembly layer has the same number of nodes. Hyperbolic tangent is used as the activation function.
@@
Nodal FML N/A $0,10^{-1}$ $1051036$
Example 2 Unconstrained PC-FML $5$ $0,10^{-1}$ $2575$
@@
Example 3 Unconstrained PC-FML $13$ 00 $44565$
@@
Table 1: Comparison of FML model parameterizations for Exs 1, 2, 3.
@@ 4.1 Example 1: 1D heat equation on incomplete non-uniform grid
To generate training data, first we solve the PDE from $N_{traj}=100$ randomized initial conditions of the form
@@
The method from Section 3.4 is used to determine an appropriate reduced basis of dimension $N_{red}=2$ for both $\sigma=0,10^{-1}$. Figure 3 (left) shows the details. As indicated by the training data generation, $b_{1}(x)=\sin(\pi x)$ and $b_{2}(x)=\sin(2\pi x)$ represent optimal basis functions, however this fact is not used in the PC-FML approach.
@@
We note the success of these methods out until $T=10$ despite having only “seen” dynamics up to $T=2$ in the training data. Table 1 shows that the failing nodal FML model uses $151036$ parameters, while the more successful PC-FML models use just $528$, $328$, and $128$ for the Unconstrained, Constrained, and Fixed options, respectively.
Figure 6: Ex. 1 (noiseless): Example trajectory (left, center) and average error (right).
Figure 7: Ex. 1 (noisy): Example trajectory (left, center) and average error (right).

### v2

@@ Abstract
We present a computational technique for modeling the evolution of dynamical systems in a reduced basis, with a focus on the challenging problem of modeling partially-observed partial differential equations (PDEs) on high-dimensional non-uniform grids. We address limitations of previous work on data-driven flow map learning in the sense that we focus on noisy and limited data to move toward data collection scenarios in real-world applications. Leveraging recent work on modeling PDEs in modal and nodal spaces, we present a neural network structure that is suitable for PDE modeling with noisy and limited data available only on a subset of the state variables or computational domain. In particular, spatial grid-point measurements are reduced using a learned linear transformation, after which the dynamics are learned in this reduced basis before being transformed back out to the nodal space. This approach yields a drastically reduced parameterization of the neural network compared with previous flow map models for nodal space learning. This allows for rapid high-resolution simulations, enabled by smaller training data sets and reduced training times.
@@ 1.1 Motivation
When modeling high-dimensional systems, e.g. systems modeled traditionally by partial differential equations (PDEs) with one or more spatial variables observed on a fine spatial grid, NN models require many parameters to account for different system behaviors. In moving from toy problems to real applications, a significant challenge throughout the field is the curse of dimensionality, which is the phenomenon that multiple system observations (training data samples) are required to accurately learn each model parameter without over-fitting. The more complex and high-dimensional a system is, the more parameters are needed in the model, and hence the more system observations are needed for training data. On the other hand, real-world data collection scenarios (e.g. storage limitations) limit the spatial grid on which measurements are made, the time at which observations are made, the state variables that are actually observed, and the amount of data collected. If only a few noisy data are available to train a highly parameterized model, accuracy and generalizability suffer. Additionally, high-resolution simulators that generate scientific data often cannot be utilized for analysis locally by different researchers. Therefore, this paper addresses the immediate need for surrogate modeling methods that require fewer parameters to enable efficient and scalable data-driven modeling of unknown high-dimensional dynamics while maintaining prediction accuracy and stability, especially when partial observations demand memory (i.e. a time history) of inputs. This paper focuses on learning flow maps in a linear reduced dimension basis to reduce the number of model parameters so that limited training data can be used productively to enable locally-executable and rapid high-resolution simulations.
@@ 1.2 Literature Review
Finally, several other papers, [1, 3, 10, 19, 26, 32], have similar goals of approximating dynamical system solutions in a reduced basis, albeit in different contexts. In particular, [1] exclusively focuses on steady-state PDEs. Citations [3, 32] focus on discovering which, and how many, components from high-dimensional observations are in fact important to system dynamics, while in what follows we focus on working with high-dimensional observations of known states. In [10], the authors use neural networks to solve a known PDE in a reduced basis. In [26], authors consider a partially-observed heat equation but observed at only 3 points throughout the domain, with little need for reduction of high-dimensional measurements and focusing more on reducing the dynamics. In [19], the authors consider a generalized Koopman framework similarly framed in the context of the Mori-Zwanzig formulation, but only consider a long ($10^{7}$ time steps), noiseless trajectory from the Lorenz ’96 ODE.
@@ 1.3 Our focus and contributions
Much like in [7], this paper considers data-driven flow map learning of PDEs from incomplete solution data, motivated by collection of high-resolution scientific data on either a partial domain or a subset of state variables, e.g. as is done in large-scale climate simulation. In other words, we consider the case of modeling an unknown PDE system using measurable data that represent partial information of the system. It is important to consider incomplete solution data because in real applications it is often difficult, if not impossible, to collect data of all the state variables of an unknown system in all areas of the relevant domain. It may even be difficult to identify all state variables in an unknown system.
Researchers performing analysis of a system on small local machines often do not have direct access to the simulators or systems themselves which are generating this type of high-dimensional scientific data. It is therefore desirable to train a surrogate model from limited data that enables approximate but inexpensive high-resolution simulations that can be executed locally. Examples in [7] used up to $10,000$ noise-less training samples to train networks with up to hundreds of thousands of parameters with high accuracy. This demonstrated the not-so-surprising result that with nearly unlimited training time and data generated from densely packed randomized initial conditions, arbitrary accuracy is possible. However, the question now moves to the limited data scenario. As training data comes at a cost in applications, it is desirable to minimize the number of parameters in the model while still capturing accurate system dynamics. In what follows, we limit data to no more than 100 samples in all examples, and add noise. This combination of factors causes the nodal memory FML models of [7] to fail, motivating this exploration.
@@
Therefore, the main contribution of the paper is a memory-based framework for flow map learning of PDEs in a learned lower-dimensional modal space from incomplete, limited, and noisy data. In particular, to reduce the measurements, we introduce unconstrained, constrained, and fixed (or pre-trained) approaches to learning common reduced bases for the training trajectories motivated by principal component analysis (PCA). We focus on linear bases to maintain small parametrizations that enable inexpensive simulation, intentionally avoiding data- and training-intensive nonlinear autoencoder-decoders. To reduce the network parameterization, we derive the existence of an ODE flow map model and additionally propose a matrix form to further reduce the network parameterization. Finally, an adjustment to the overall ResNet structure radically improves the FML strategy’s robustness to noise. In a set of both linear and nonlinear numerical examples, we compare these options against each other as well as the existing nodal model from [7] to demonstrate that our method is indeed highly effective and accurate.
@@ 3.2.1 Reduced basis for nodal measurements
$\delta_{n}^{(l)}=\left\|\mathbf{V}_{n}^{(l)}-\tilde{\mathbf{V}}^{(l)}_{n}\right\|^{2}_{2},$ (21)
for $l=1,\ldots,N_{traj}$ and $n=1,\ldots,N_{mem}+N_{rec}$ as in (17). Generally, if $N_{red}$ were fixed, we might seek an optimal basis that accurately approximates training data by minimizing over $\delta_{n}^{(l)}$ for all trajectories and time steps in the training data. We note that an optimal basis may be significantly different from those generating each training instance’s exact finite-term series representation of the finite-dimensional grid measurements.
@@
A focus of the ensuing technique is on finding matrices $\Pi_{in}$ and $\Pi_{out}$, that balance minimization of the error $\delta_{n}^{(l)}$ and the reduced dimension $N_{red}$ over the entire training dataset (17). As described in Section 3.5, several approaches are taken, motivated by principal component analysis (PCA), also known as principal orthogonal decomposition (POD). E.g., perhaps the most straightforward approach is to use fixed transformations based on the PCA weight matrix which can be pre-computed from the training data. This would enforce an accurate approximation of $\mathbf{V}_{n}$ by $\Pi_{out}\Pi_{in}\mathbf{V}_{n}$ along with $\Pi_{in}=\Pi_{out}^{T}$ and $\Pi_{out}$ having orthogonal columns.
@@ 3.2.2 Partial reduced dynamics
the solution of which has an approximate correspondence to the solution of the PDE in $\mathbf{V}$. This permits learning the evolution of the system in $\mathbb{R}^{N_{red}}$ instead of $\mathbb{R}^{N_{full}}$. When the governing PDE is unknown, the system for $\hat{\mathbf{V}}$ is unknown as well. Since $\mathbf{V}$ represent partial observations of the entire PDE system, $\hat{\mathbf{V}}$ represent partial observations of some entire system, say for $\hat{\mathbf{U}}$, the modal coefficients of some approximation of the full system.
@@
$\bar{\mathbf{V}}_{n+1}=\Pi_{out}\Pi_{in}\bar{\mathbf{V}}_{n}+\Pi_{out}\mathcal{M}\left(\Pi_{in}\bar{\mathbf{V}}_{n},\ldots,\Pi_{in}\bar{\mathbf{V}}_{n-(N_{mem}-1)}\right),$ (28)
where $\mathcal{M}$ approximates the modal space coefficient flow map, and the bar notation $\bar{\mathbf{V}}$ denotes that model inputs and outputs are approximations of $\mathbf{V}$. After extensive numerical experimentation, we find that the transformation of $\bar{\mathbf{V}}_{n}$ into and out of the reduced space is essential to dealing with noisy nodal measurements. Essentially, $\Pi_{out}\Pi_{in}$ functions as a denoiser. If measurements are noise-free, it can be removed.
@@ 3.4 Determining reduced order dimension
(a) Ex. 1: Heat
(b) Ex. 2: Burgers’
(c) Ex. 3: Wave 1D
(d) Ex. 4: Shallow Water
(e) Ex. 5: Wave 2D
(f) Ex. 6: Navier-Stokes
Figure 3: Determining reduced order dimension.
Figure 3 shows the decay of the singular values in each of the six forthcoming numerical experiments with and without noise. We see, e.g., in Ex. 5 the noiseless singular values of the training data drop off gradually after $N_{red}=13$, while noise is dominant after $N_{red}=5$ in the noisy observations. Additionally, this SVD computation exposes potential basis elements $b_{j}(\mathbf{x})$ for $j=1,\ldots,N_{red}$ in the columns of $V_{red}$, which can be used to explain the decomposition and interpret the dynamics. Figures 4 and 5 show the 13- and 5-dimensional bases for the noiseless and noisy cases, respectively, in Example 3. We see similar shapes for the first 5 elements, while we also notice noise creeping into the $b_{4}(\mathbf{x})$ and $b_{5}(\mathbf{x})$ in the noisy measurements. This demonstrates why after this point, including extra basis elements is not advantageous as they will have a large noise component that may actually be detrimental to reconstructing the unknown dynamics.
Figure 4: Ex. 5 (noiseless 2D wave): Reduced basis elements $b_{j}(\mathbf{x})$ for Fixed PC-FML.
Figure 5: Ex. 5 (noisy 2D wave): Reduced basis elements $b_{j}(\mathbf{x})$ for Fixed PC-FML.
@@ 3.5 Model Training
Essentially, the options correspond to the amount of freedom we give the data and the network to learn the transformation. Indeed the inspiration for the general method was the Fixed regime, which can be relaxed to the Constrained regime’s orthogonality and symmetry or completely dropped in the Unconstrained regime. [^6] We admit that, despite theoretical guarantees of the SVD, the network may find a better FML basis! However, as stated in examples below, as $N_{red}$ increases, such as in nonlinear PDEs, the Constrained and Unconstrained approaches struggle to simultaneously learn appropriate reduced bases and system dynamics. Considering they are only trying to learn linear bases, this provides further support for the avoidance of learning highly-parameterized nonlinear bases, e.g. via autoencoder-decoders. Notably, in all cases the loss is measured in the nodal domain. This is significant, and differs from [4, 31], since the network parameters are trained to minimize loss in the nodal domain. If instead training is only executed in the modal domain, expanding back out could cause larger errors in the nodal domain, which is the focus.
@@ 4 Computational Studies
  • Time-step: $\Delta t=10^{-2}$ apart from Ex. 4 which uses $\Delta t=8\times 10^{-3}$.
@@
  • Recurrent loss: $N_{rec}=10$ steps for linear PDEs, and $N_{rec}=20$ for nonlinear.
@@
  • Noise: additive zero-mean Gaussian noise is considered at two noise levels, with standard deviations $\sigma=0$ and $\sigma=10^{-1}$ or $10^{-2}$, referred to generically as noiseless and noisy.
  • PC-FML network: networks $\mathbf{M}$ have $3$ hidden layers with up to $60$ nodes per layer determined by the multiple of $10$ above the reduced dimension. E.g., in Ex. 2 we have $N_{red}=14$ and so $20$ nodes are used in each layer. Hyperbolic tangent is used as the activation function.
  • Nodal FML network size: comparison networks from [7] have $5$ disassembly channels with $1$ hidden layer with either $N_{grid}$ or $N_{red}$ nodes per layer, where using $N_{red}$ represents a naive attempt at dimensionality reduction. The assembly layer has $5$ nodes. Hyperbolic tangent is used as the activation function.
@@
Nodal FML N/A $0,10^{-1}$ $151036$
Example 2 (Burgers’) Unconstrained PC-FML $14$ 00 $15154$
Constrained PC-FML $14$ 00 $10954$
Fixed PC-FML $14$ 00 $6754$
Nodal FML N/A 00 $442606$
Unconstrained PC-FML $6$ $10^{-1}$ $5096$
Constrained PC-FML $6$ $10^{-1}$ $3296$
Fixed PC-FML $6$ $10^{-1}$ $1496$
Nodal FML N/A $10^{-1}$ $190566$
Example 3 (Wave 1D) Unconstrained PC-FML $5$ $0,10^{-1}$ $2575$
@@
Example 4 (Shallow Water) Fixed PC-FML $52$ 00 $93752$
Nodal FML N/A 00 $1093296$
Fixed PC-FML $6$ $10^{-2}$ $9891$
Nodal FML N/A $10^{-2}$ $232091$
Example 5 (Wave 2D) Unconstrained PC-FML $13$ 00 $44565$
@@
Example 6 (Navier-Stokes) Fixed PC-FML $46$ 00 $53496$
Nodal FML N/A 00 $12087766$
Fixed PC-FML $26$ $10^{-2}$ $18296$
Nodal FML N/A $10^{-2}$ $6837666$
Table 1: Comparison of FML model parametrizations.
@@ 4.1 Example 1: 1D heat equation on incomplete non-uniform grid
To generate training data, first we solve the PDE numerically from $N_{traj}=100$ randomized initial conditions of the form
@@
The method from Section 3.4 is used to determine an appropriate reduced basis of dimension $N_{red}=2$ for both $\sigma=0,10^{-1}$. Figure 3 shows the details. As indicated by the training data generation, $b_{1}(x)=\sin(\pi x)$ and $b_{2}(x)=\sin(2\pi x)$ represent optimal basis functions, however this fact is not used in the PC-FML approach.
@@
We note the success of these methods out until $T=10$ despite having only “seen” dynamics up to $T=2$ in the training data. Table 1 shows that the failing nodal FML model uses $151036$ parameters, while the more successful PC-FML models use just $1052$, $852$, and $652$ for the Unconstrained, Constrained, and Fixed options, respectively.
Figure 6: Ex. 1 (noiseless 1D heat): Example trajectory (left, center) and average error (right).
Figure 7: Ex. 1 (noisy 1D heat): Example trajectory (left, center) and average error (right).
4.2 Example 2: 1D Burgers’ equation on incomplete uniform grid
We next consider the nonlinear Burgers’ equation with viscosity
$u_{t}+uu_{x}$ $=0.05u_{xx}$ (38)
on the domain $\Omega=[-\pi,\pi]$ with periodic boundary conditions. The overall behavior of the PDE is characterized by formation of a shock but then dissipation before a discontinuity forms.
To generate training data, first we numerically solve the PDE from $N_{traj}=100$ randomized initial conditions as in (37) where $\alpha_{1},\alpha_{2}\sim U[-1,1]$, until $T=3$ ($300$ observed time steps). One chunk of length $N_{mem}+N_{rec}$ is selected randomly from each trajectory to form the training data set. The solutions are observed on a uniform grid of $N_{grid}=300$ points representing roughly the middle $60\%$ of the domain, $[-3\pi/5,3\pi/5]\subset\Omega$.
The method from Section 3.4 is used to determine an appropriate reduced basis of dimension $N_{red}=14$ for $\sigma=0$ and $N_{red}=6$ for $10^{-1}$. Figure 3 shows the details.
Validation testing also uses $100$ trajectories drawn similarly to (37). On the left and center of Figures 8 (noiseless) and 9 (noisy), an example test trajectory is shown at various prediction times, and on the right the average absolute $\ell_{2}$ error the test trajectories is shown over $300$ time steps. In the noiseless case, the Fixed PC-FML strategy performs best. We hypothesize that because the $N_{red}=14$ is relatively high (representing a $N_{red}N_{mem}=280$ element input to $\mathbf{M}$), that Constrained the Unconstrained PC-FML have the additional challenge of discovering a basis rather than having one provided. Interestingly, all the PC-FML strategies are successful in the noisy case at learning denoised evolution of the system. Table 1 shows that the nodal FML models use between 29 and 127 times more parameters compared with PC-FML.
Figure 8: Ex. 2 (noiseless 1D Burgers’): Example trajectory (left, center) and average error (right).
Figure 9: Ex. 2 (noisy 1D Burgers’): Example trajectory (left, center) and average error (right).